# SoftHGNN: Soft Hypergraph Neural Networks for General Visual Recognition

### v1

@@ V-D1 Image Classification
In the image classification task, we adopt the classical Vision Transformer (ViT) [5] as our baseline. As illustrated in Figure 2, we incrementally introduce different modules at the end of the ViT encoder for comparison: Self-Attention [5], HGNN [21], the proposed SoftHGNN, and SoftHGNN-SeS (with the Sparse Hyperedge Selection). Table I presents the results on the CIFAR-10 [16] and CIFAR-100 [16] datasets. From the table, It can be observed that adding an additional Self-Attention module brings no significant improvement, as ViT itself already relies on attention mechanism. This further highlights the inherent limitation of self-attention in modeling only explicit, pairwise visual associations. When introducing the HGNN module, performance slightly drops on CIFAR-10 and slightly improves on CIFAR-100. This instability is attributed to the hard binary nature of HGNN’s hyperedge assignments. In our experience, HGNNs often require extensive and delicate hyperparameter tuning to achieve competitive performance. In contrast, when the proposed SoftHGNN module is integrated, we observe a significant improvement of 1.9% on CIFAR-10 and 1.8% on CIFAR-100. Moreover, when we increase the number of soft hyperedges and apply sparse hyperedge selection along with the load-balancing regularization, the model achieves further gains. Specifically, SoftHGNN-SeS improves upon the baseline by 2.5% and 2.4% on CIFAR-10 and CIFAR-100, respectively.
@@
V-D2 Crowd Counting
In the crowd counting task, we adopt the Transformer-based CCTrans [35] as our baseline, and introduce Self-Attention [5], HGNN [21], SoftHGNN, and SoftHGNN-SeS into the multi-scale feature fusion stage, as illustrated in Figure 2. Table II presents the experimental results on the ShanghaiTech Part-A and Part-B datasets. We the methods using two different backbone networks: one is the Twins-PCPVT [9] used by the original CCTrans, and the other is the more lightweight and efficient Pyramid Vision Transformer (PVT) [6]. Under both backbones, the proposed SoftHGNN series significantly outperform both Self-Attention and the traditional HGNN. When using Twins-PCPVT as the backbone, SoftHGNN series achieve the largest improvements of 3.6% in MAE and 7.9% in MSE on Part-A, and 7.9% in MAE and 13.1% in MSE on Part-B. When using PVT as the backbone, the improvements brought by the SoftHGNN series are even more substantial: the MAE and MSE improve by up to 3.7% and 7.8% on Part-A, and by up to 10.8% and 16.0% on Part-B. Furthermore, it is worth noting that even with significant improvements, SoftHGNN series still maintains the efficiency of parameters. Specifically, SoftHGNN introduces only about 1M additional parameters, while SoftHGNN-SeS adds approximately 6M. This demonstrates that our proposed SoftHGNN series effectively capture high-order semantic associations that are overlooked by traditional HGNN and Self-Attention, thereby enhancing the model’s representational capacity in an efficient manner.
TABLE III: Quantitative results of nano size model object detection performance on MS COCO dataset.
Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param.
Gold-YOLO-N 55.7 39.6 12.1 G 2.92 ms 5.6 M
Hyper-YOLO-T 54.5 38.5 9.6 G 2.5 ms 3.1 M
YOLOv8-N 52.6 37.4 8.7 G 1.77 ms 3.2 M
YOLOv9-T 53.1 38.3 7.7 G 2.48 ms 2.0 M
YOLOv10-N 53.8 38.5 6.7 G 1.84 ms 2.3 M
YOLO11-N 55.3 39.4 6.5 G 1.50 ms 2.6 M
YOLO11-N 53.6 38.2 6.5 G 1.59 ms 2.6 M
+ SoftHGNN-N 55.8 ($\uparrow$2.2) 40.2 ($\uparrow$2.0) 8.0 G 1.99 ms 3.2 M
YOLOv12-N 55.4 40.1 6.0 G 1.60 ms 2.5 M
YOLOv12-N 55.1 39.6 6.0 G 1.63 ms 2.5 M
+ SoftHGNN-N 57.0 ($\uparrow$1.9) 41.3 ($\uparrow$1.7) 7.4 G 1.99 ms 3.1 M
TABLE IV: Quantitative results of small size model object detection performance on MS COCO dataset.
Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param.
Gold-YOLO-S 62.5 45.4 46.0 G 3.82 ms 21.5 M
Hyper-YOLO-S 65.1 48.0 39.0 G 4.7 ms 14.8 M
YOLOv8-S 61.8 45.0 28.6 G 2.33 ms 11.2 M
YOLOv9-S 63.4 46.8 26.4 G 3.46 ms 7.1 M
YOLOv10-S 63.0 46.3 21.6 G 2.49 ms 7.2 M
YOLO11-S 63.9 46.9 21.5 G 2.50 ms 9.4 M
YOLO11-S 62.2 45.6 21.5 G 2.57 ms 9.4 M
+ SoftHGNN-S 63.8 ($\uparrow$1.6) 47.1 ($\uparrow$1.5) 27.4 G 3.16 ms 12.1 M
YOLOv12-S 64.1 47.6 19.4 G 2.42 ms 9.1 M
YOLOv12-S 63.8 47.0 19.4 G 2.49 ms 9.1 M
+ SoftHGNN-S 64.8 ($\uparrow$1.0) 48.0 ($\uparrow$1.0) 25.2 G 3.17 ms 11.4 M
TABLE V: Quantitative results of medium size model object detection performance on MS COCO dataset.
Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param.
Gold-YOLO-M 67.0 49.8 87.5 G 6.38 ms 41.3 M
Hyper-YOLO-M 69.0 52.0 103.3 G 9.0 ms 33.3 M
YOLOv8-M 67.2 50.3 78.9 G 5.09 ms 25.9 M
YOLOv9-M 68.1 51.4 76.3 G 6.18 ms 20.0 M
YOLOv10-M 68.1 51.1 59.1 G 4.74 ms 15.4 M
YOLO11-M 68.5 51.5 68.0 G 4.70 ms 20.1 M
YOLO11-M 67.3 50.5 68.0 G 5.04 ms 20.1 M
+ SoftHGNN-M 68.1 ($\uparrow$0.8) 51.1 ($\uparrow$0.6) 85.0 G 6.07 ms 25.1 M
YOLOv12-M 69.5 52.6 59.8 G 4.27 ms 19.6 M
YOLOv12-M 68.6 51.9 59.8 G 4.75 ms 19.6 M
+ SoftHGNN-M 69.2 ($\uparrow$0.6) 52.2 ($\uparrow$0.3) 80.2 G 6.01 ms 26.3 M
V-D3 Object Detection
@@
F2SoftHG first fuses multi-scale features using a convolutional block. Then, the fused feature map is fed into a group of SoftHGNNs to model high-order semantic associations. Finally, the output of F2SoftHG is redistributed to each level and fused with the feature maps of the corresponding levels via concatenation and convolution. In this way, F2SoftHG fully leverages the advantages of both SoftHGNN and convolution to realize high-order semantic association extraction across spatial and scale dimensions, promoting multi-object and object-background interactions at the global level.
We conduct experiments on the Nano, Small, and Medium variants of YOLO11 and YOLOv12, as shown in Table III, Table IV, and Table V. For the Nano scale, our method improves $\text{AP}_{50}^{val}$ by 2.2% and 1.9%, and $\text{AP}_{50:95}^{val}$ by 2.0% and 1.7% over YOLO11 and YOLOv12, respectively. For the Small scale, it improves $\text{AP}_{50}^{val}$ by 1.6% and 1.0%, and $\text{AP}_{50:95}^{val}$ by 1.5% and 1.0%. For the Medium scale, it achieves an increase of 0.8% and 0.6% in $\text{AP}_{50}^{val}$, and 0.6% and 0.3% in $\text{AP}_{50:95}^{val}$ compared to YOLO11 and YOLOv12, respectively. In addition, our method introduces only a small number of additional parameters and computational cost. Compared with YOLOv12-S, our method introduces only an additional 2.3M parameters and 5.8G FLOPs, while still maintaining real-time performance and high computational efficiency.
@@ V-E1 Hyperparameter Sensitivity
Next, we fix the number of heads to 8 and maintain all other settings unchanged, while varying the number of soft hyperedges (4, 8, 16, 32). The results show that SoftHGNN performs better when the number of soft hyperedges is set to 8 or 16, with the best MAE and MSE achieved at 8 soft hyperedges. It is noteworthy that in SoftHGNN, increasing the number of soft hyperedges does not always lead to better performance, despite providing a larger capacity for learnable parameters. We attribute this to two reasons. On the one hand, the number of effective high-order semantic relations is directly related to the complexity of the scenes in the dataset. In most cases, a relatively small number of soft hyperedges (such as 8 or 16) is sufficient to capture enough high-order semantic relations. On the other hand, when the number of soft hyperedges becomes too large, the model tends to capture less important or redundant relations, or overfit to specific bad-case associations. SoftHGNN-SeS is exactly proposed to solve this issue. As shown in Table I and Table II, SoftHGNN-SeS extends the learning capacity of SoftHGNN by first providing a larger number of dynamic candidate soft hyperedges, and then sparsely selecting only the most important ones. This approach allows the model to benefit from a larger hyperedge space while avoiding redundancy and overfitting, thus significantly enhancing the upper bound of its learning ability.
V-E2 Computational Efficiency

### v2

@@ 5.4.1 Image Classification
In the image classification task, we adopt the classical Vision Transformer (ViT) [4] as our baseline. As illustrated in Figure 2, we incrementally introduce different modules at the end of the ViT encoder for comparison: Self-Attention [4], HGNN [22], the proposed SoftHGNN, and SoftHGNN-SeS (with the Sparse Hyperedge Selection). Table 1 presents the results on the CIFAR-10 [16] and CIFAR-100 [16] datasets. From the table, It can be observed that adding an additional Self-Attention module brings no significant improvement, as ViT itself already relies on attention mechanism. This further highlights the inherent limitation of self-attention in modeling only explicit, pairwise visual associations. When introducing the HGNN module, performance slightly drops on CIFAR-10 and slightly improves on CIFAR-100. This instability is attributed to the hard binary nature of HGNN’s hyperedge assignments. In our experience, HGNNs often require extensive and delicate hyperparameter tuning to achieve competitive performance. In contrast, when the proposed SoftHGNN module is integrated, we observe a significant improvement of 1.9% on CIFAR-10 and 1.8% on CIFAR-100. Moreover, when we increase the number of soft hyperedges and apply sparse hyperedge selection along with the load-balancing regularization, the model achieves further gains. Specifically, SoftHGNN-SeS improves upon the baseline by 2.5% and 2.4% on CIFAR-10 and CIFAR-100, respectively.
Table 4: Quantitative results of small size model object detection performance on MS COCO dataset.
Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param.
Gold-YOLO-S [59] 62.5 45.4 46.0 G 3.82 ms 21.5 M
Hyper-YOLO-S [15] 65.1 48.0 39.0 G 4.7 ms 14.8 M
YOLOv8-S [60] 61.8 45.0 28.6 G 2.33 ms 11.2 M
YOLOv9-S [61] 63.4 46.8 26.4 G 3.46 ms 7.1 M
YOLOv10-S [62] 63.0 46.3 21.6 G 2.49 ms 7.2 M
YOLO11-S [57] 63.9 46.9 21.5 G 2.50 ms 9.4 M
YOLO11-S [57] 62.2 45.6 21.5 G 2.57 ms 9.4 M
YOLO11-SoftHGNN-S 63.8 ($\uparrow$1.6) 47.1 ($\uparrow$1.5) 27.4 G 3.16 ms 12.1 M
YOLOv12-S [58] 64.1 47.6 19.4 G 2.42 ms 9.1 M
YOLOv12-S [58] 63.8 47.0 19.4 G 2.49 ms 9.1 M
YOLOv12-SoftHGNN-S 64.8 ($\uparrow$1.0) 48.0 ($\uparrow$1.0) 25.2 G 3.17 ms 11.4 M
Table 5: Quantitative results of medium size model object detection performance on MS COCO dataset.
Method $\text{AP}_{50}^{val}$ $\text{AP}_{50:95}^{val}$ FLOPs Latency Param.
Gold-YOLO-M [59] 67.0 49.8 87.5 G 6.38 ms 41.3 M
Hyper-YOLO-M [15] 69.0 52.0 103.3 G 9.0 ms 33.3 M
YOLOv8-M [60] 67.2 50.3 78.9 G 5.09 ms 25.9 M
YOLOv9-M [61] 68.1 51.4 76.3 G 6.18 ms 20.0 M
YOLOv10-M [62] 68.1 51.1 59.1 G 4.74 ms 15.4 M
YOLO11-M [57] 68.5 51.5 68.0 G 4.70 ms 20.1 M
YOLO11-M [57] 67.3 50.5 68.0 G 5.04 ms 20.1 M
YOLO11-SoftHGNN-M 68.1 ($\uparrow$0.8) 51.1 ($\uparrow$0.6) 85.0 G 6.07 ms 25.1 M
YOLOv12-M [58] 69.5 52.6 59.8 G 4.27 ms 19.6 M
YOLOv12-M [58] 68.6 51.9 59.8 G 4.75 ms 19.6 M
YOLOv12-SoftHGNN-M 69.2 ($\uparrow$0.6) 52.2 ($\uparrow$0.3) 80.2 G 6.01 ms 26.3 M
5.4.2 Crowd Counting
In the crowd counting task, we adopt the Transformer-based CCTrans [39] as our baseline, and introduce Self-Attention [4], HGNN [22], SoftHGNN, and SoftHGNN-SeS into the multi-scale feature fusion stage, as illustrated in Figure 2. Table 2 presents the experimental results on the ShanghaiTech Part-A and Part-B datasets. We the methods using two different backbone networks: one is the Twins-PCPVT [54] used by the original CCTrans, and the other is the more lightweight and efficient Pyramid Vision Transformer (PVT) [5]. Under both backbones, the proposed SoftHGNN series significantly outperform both Self-Attention and the traditional HGNN. When using Twins-PCPVT as the backbone, SoftHGNN series achieve the largest improvements of 3.6% in MAE and 7.9% in MSE on Part-A, and 7.9% in MAE and 13.1% in MSE on Part-B. When using PVT as the backbone, the improvements brought by the SoftHGNN series are even more substantial: the MAE and MSE improve by up to 3.7% and 7.8% on Part-A, and by up to 10.8% and 16.0% on Part-B. Furthermore, it is worth noting that even with significant improvements, SoftHGNN series still maintains the efficiency of parameters. Specifically, SoftHGNN introduces only about 1M additional parameters, while SoftHGNN-SeS adds approximately 6M. This demonstrates that our proposed SoftHGNN series effectively capture high-order semantic associations that are overlooked by traditional HGNN and Self-Attention, thereby enhancing the model’s representational capacity in an efficient manner.
@@ 5.4.2 Crowd Counting
5.4.3 Object Detection
@@
F2SoftHG first fuses multi-scale features using a convolutional block. Then, the fused feature map is fed into a group of SoftHGNNs to model high-order semantic associations. Finally, the output of F2SoftHG is redistributed to each level and fused with the feature maps of the corresponding levels via concatenation and convolution. In this way, F2SoftHG fully leverages the advantages of both SoftHGNN and convolution to realize high-order semantic association extraction across spatial and scale dimensions, promoting multi-object and object-background interactions at global level.
We conduct experiments on the Nano, Small, and Medium variants of YOLO11 and YOLOv12, as shown in Table 3, Table 4, and Table 5. For the Nano scale, our method improves $\text{AP}_{50}^{val}$ by 2.2% and 1.9%, and $\text{AP}_{50:95}^{val}$ by 2.0% and 1.7% over YOLO11 and YOLOv12, respectively. For the Small scale, it improves $\text{AP}_{50}^{val}$ by 1.6% and 1.0%, and $\text{AP}_{50:95}^{val}$ by 1.5% and 1.0%. For the Medium scale, it achieves an increase of 0.8% and 0.6% in $\text{AP}_{50}^{val}$, and 0.6% and 0.3% in $\text{AP}_{50:95}^{val}$ compared to YOLO11 and YOLOv12, respectively. In addition, our method introduces only a small number of additional parameters and computational cost. Compared with YOLOv12-S, our method introduces only an additional 2.3M parameters and 5.8G FLOPs, while still maintaining real-time performance and high computational efficiency.
@@ 5.5.1 Hyperparameter Sensitivity
Next, we fix the number of heads to 8 and maintain all other settings unchanged, while varying the number of soft hyperedges (4, 8, 16, 32). The results show that SoftHGNN performs better when the number of soft hyperedges is set to 8 or 16, with the best MAE and MSE achieved at 8 soft hyperedges. It is noteworthy that in SoftHGNN, increasing the number of soft hyperedges does not always lead to better performance, despite providing a larger capacity for learnable parameters. We attribute this to two reasons. On the one hand, the number of effective high-order semantic relations is directly related to the complexity of the scenes in the dataset. In most cases, a relatively small number of soft hyperedges (such as 8 or 16) is sufficient to capture enough high-order semantic relations. On the other hand, when the number of soft hyperedges becomes too large, the model tends to capture less important or redundant relations, or overfit to specific bad-case associations. SoftHGNN-SeS is exactly proposed to solve this issue. As shown in Table 1 and Table 2, SoftHGNN-SeS extends the learning capacity of SoftHGNN by first providing a larger number of dynamic candidate soft hyperedges, and then sparsely selecting only the most important ones. This approach allows the model to benefit from a larger hyperedge space while avoiding redundancy and overfitting, thus significantly enhancing the upper bound of its learning ability.
5.5.2 Computational Efficiency
@@ 6 Conclusion
7 Data Availability
The datasets used in this study are available in the following public domain resources: